OpenStudy (anonymous):

A system of two linear equations has an infinite number of solutions. How is this possible?

OpenStudy (anonymous):

when the linear equations represent the same line, then it has infinite number

OpenStudy (anonymous):

If they are consistent, example: x + y = 4 2x + 2y = 8

OpenStudy (anonymous):

ohh tanks

OpenStudy (anonymous):

3x – 4y = 25 3x + 2y = 1 could you help on this one?

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Subtract the two equations, you get -6y = 24

OpenStudy (anonymous):

y = -4

OpenStudy (anonymous):

3, -4?

OpenStudy (anonymous):

yes